OpenStudy (anonymous):

1) You can find the distance between two points on a number line by using the Pythagorean theorem. True False 2) The distance between two points on a coordinate plane can always be found by subtracting their x- or y-coordinates. True False 3) Sonja's house is 4 blocks west and 1 block south of the center of town. Her school is 3 blocks east and 2 blocks north of the center of town. What is the direct distance from Sonja's house to her school? (Hint: The center of town should be the origin, and north is up.) 10 blocks 4.5 blocks 7.1 blocks 7.6 blocks

OpenStudy (nirmalnema):

1) false we have to use distance formula 2) true 3) \[\sqrt{(x1-x1)^{2}+(y1-y2)^{2}}\] = rt ov (3+4)^2 + (2+1)^2 =rt ov 49+9 =rt ov 58 = 7.6